OpenStudy (anonymous):

algebra 1 really difficult medal shall be rewarded!

OpenStudy (anonymous):

@AnElephant

OpenStudy (anonymous):

6. -2r*8r+-2r*5 7. 4a*a^2+4a*7a+4*4a now just simplify them

OpenStudy (anonymous):

be sure to include your 'r's otherwise the problem won't be right

OpenStudy (anonymous):

6) -16 - 10

OpenStudy (anonymous):

I think -16r -10r <--

OpenStudy (anonymous):

Yes, now for 7?

OpenStudy (anonymous):

-16r^2 -10r loveyou*69

OpenStudy (anonymous):

4a*a^2= 4a^3 4a*7a = 28a^2 4*4a = 16a

OpenStudy (anonymous):

7)4a(a^2 + 7a + 4) 4a • a^2 = 4a^3 4a • 7a = 28^2 4a • 4 = 16 a 4a^3 + 28a^2 + 16a

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

Anytime, glad to help. (:

OpenStudy (anonymous):

hy I am going to give you a medal @AnElephant can you give loveyou*69 a medal too please?

OpenStudy (anonymous):

thanks!